OpenStudy (anonymous):

taking the derivative of x^(2/3) using lim f(x+h)-f(x)/h

OpenStudy (anonymous):

Please I Really Need Help[ With This Problem

OpenStudy (anonymous):

it would help if you got rid of the rational exponent and wrote \[\lim_{h\rightarrow 0}\frac{\sqrt[3]{(x+h)^2}-\sqrt[3]{x^2}}{h}\]

OpenStudy (anonymous):

damn i just wrote a beginning and it got lost!

OpenStudy (anonymous):

oh there it is!

OpenStudy (anonymous):

well that way than

OpenStudy (anonymous):

can you show the rest of the steps to solving the problem?

OpenStudy (anonymous):

satellite?